# FeatSharp: Your Vision Model Features, Sharper

### v1

@@ 4.3 Semantic Segmentation
Semantic segmentation has the potential to benefit from increased resolution, as it allows for label contours to be more precise, and potentially for regions to be recovered that are otherwise too small. The first setting we evaluate on is we train both FeatUp and FeatSharp at 2×2$\times$2 × and 4×4$\times$4 × upsampling, both using PHI-S. We resize the input size to be the featurizer’s native input resolution, which we call “1×1$\times$1 × Input Size”, and we also consider “2×2$\times$2 × Input Size”, where we double the input size, and feed directly to the featurizer in the case of “Baseline”, or we allow the upsampler to have higher resolution guidance while keeping the featurizer input fixed at 1×1$\times$1 × resolution. We show these results in figure 7. In most cases, both upsampling algorithms produce higher quality segmentations than the baseline, however, FeatUp is worse than the “Baseline 2×2$\times$2 ×” method for RADIOv2.5-L and ViT. In all cases, FeatSharp is superior to both FeatUp and also the baselines by significant margins. We even improve upon SOTA RADIO’s published result of 51.47 with a 2×2$\times$2 × upsampling combined with 2×2$\times$2 × input size, producing a model that attains 53.13 mIoU, a $+1.66$ mIoU improvement. RADIO itself improves with the 2×2$\times$2 × input size, but not to the same degree as with FeatSharp, with FeatSharp being $57\%$ faster. We also notice that 3×3$\times$3 × upsampling is generally worse than 2×2$\times$2 × or 4×4$\times$4 × for both upsamplers.
@@ 4.4 Agglomerative Models
In figure 8 we qualitatively visualize the DFN CLIP adaptor features learned by the radio model. We can see that each upsampling method has a substantial impact on the resulting feature maps. The baseline method exhibits strong high-frequency artifacting starting at 768px. This is likely when RADIO “mode switches” to high-resolution, which is something that (Heinrich et al., 2024) addressed for the backbone features, but apparently still exhibit for the adaptor features. We observe that Tiling and S2 exhibit not only high-frequency noise patterns like the baseline, but also obvious grid patterns, arising from the use of tiles. FeatUp appears to mode switch starting at 768px into a smooth, but low-detail feature space. FeatSharp remains smooth and highly detailed as resolution increases, however, visually, it’s still possible that the features are mode switching. We further study this tiling issue in appendix F.
Along with improvements in the adaptors, we also study the effects on the backbone features for the RADIO model. Following (Maninis et al., 2019; Lu et al., 2024) we report the MTL Gain ($\Delta_{m}$) across a suite of tasks. Unlike the prior works, instead of leveraging a single-task baseline, we instead opt to report the change relative to the baseline training run.
@@
We show the MTL Gain results in table 1. Given that the results are relative to our baseline run, S2 and FeatSharp are the only two methods to improve, however, only FeatSharp was categorically better, leading to a +0.39% improvement across all benchmarks on average. We also see that our version of RADIO with FeatSharp teachers generally does better than RADIO-AMP-L (Heinrich et al., 2024), which is the current state of the art, where we improve over it on everything except for the VILA task. We report all of the raw benchmark scores in tables 2, 3, 4 and 5 in the appendix.
@@ 5 Conclusion
We have presented a novel feature upsampling technique named FeatSharp that achieves higher multi-view fidelity than the current best method, FeatUp. We achieve this by joining FeatUp’s JBU upsampler with a mosaic of tiles, and then process with a single local attention block. We demonstrate its effectiveness on ADE20K semantic segmentation linear probing, where the use of FeatSharp improves over both baseline and FeatUp, even with the strongest segmenter, RADIO, which itself can handle hi-res inputs robustly. We then demonstrate the effectiveness of FeatSharp by employing it directly within RADIO training, enabling low-res-only teacher models to have hi-res distillation targets. In doing this, our FeatSharp-RADIO largely improves on dense vision task benchmarks, and yields an overall improvement over our reproduction baseline, which itself improves over RADIO-AMP-L, the current state of the art. We believe this work can be useful both as a drop-in extension of existing vision systems which rely on pretrained vision encoders, as well as the newly trained FeatSharp-RADIO model with hi-res teachers, which can emulate these same models. Owing to FeatSharp-RADIO’s emulation abilities, it allows us to estimate these teacher models at arbitrary resolutions, not just integer upsampling factors as restricted in FeatSharp/FeatUp’s core training algorithm. Further, combining RADIO’s “ViTDet” (Li et al., 2022) mode with these hi-res teacher emulations allows us to achieve hi-res feature maps without fully paying the quadratic penalty in number of tokens as required by standard ViTs.
@@ A.1 Benchmark Results
Table 3: Dense and Probe3D (El Banani et al., 2024) metrics. *We report numbers for evaluation at 512px, which are found in Table A5 in RADIO-AMP (Heinrich et al., 2024).
@@
Table 4: Pascal Context and NYUDv2 multitask learning metrics. Following the setup of MLoRE (Yang et al., 2024b) and RADIO-AMP (Heinrich et al., 2024) with a convolutional probe. NOTE: We’re only using their harness with a conv probe, and not using their architecture.
@@ A.2 Additional Qualitative Visualizations
In figures 9 and 10, we show more PCA feature visualizations coming from our trained RADIO models. In figure 9 we can see that RADIO learned to mimic how tiling lacks global context, as the background-only tiles use a different feature space than those with background+content.
Figure 9: Visualization of our trained RADIO’s DFN CLIP adaptor, using different teacher upsampling techniques.
@@ B.2 FeatSharp Architecture
Based on figures 1 and 2, there are important degrees of freedom in the design of the system. We demonstrate in table 6 that including the de-bias module always improves the distribution matching fidelity. Here, we look at some of the other design choices:
@@
We show the results of this ablation in table 7 for both our noisiest featurizer (SigLIP), and our cleanest (RADIO) as a sanity check that we aren’t overfitting to a particular featurizer.
@@
Bilinear 1.423 0.141 0.114 3.952 0.076 0.064
FeatUp 1.380 0.025 0.046 3.324 0.034 0.042
Tiles 1.245 0.938 0.211 2.636 0.483 0.096
@@
Bilinear + Tiles 1.451 0.229 0.127 4.216 0.141 0.073
FeatUp + Tiles 1.424 0.172 0.071 3.821 0.184 0.061
Tiles + Bilinear 1.396 0.683 0.188 3.744 0.383 0.096
Tiles + FeatUp 1.357 0.790 0.190 3.517 0.424 0.093
@@
Bilinear First 1.450 0.222 0.126 4.184 0.126 0.070
FeatUp First 1.438 0.142 0.074 3.950 0.161 0.063
Tiles First 1.395 0.680 0.183 3.756 0.396 0.095
Table 7: Ablation over different FeatSharp-2x configurations. Single Input means that we only supply the respective buffer to the FeatSharp module. For “Two Inputs“, we compare different low-res upsamplers in conjunction with tiling, and also the residual pathway, where the first value indicates the residual path. The FeatSharp module must integrate the information from the other value into the residual. “Three Inputs” is similar to Two, except that we only care about which buffer is the residual path, owing to the fact that there’s no intrinsic order preference in the weights for the secondary buffer(s). “TV Loss” stands for Total Variation Loss (Rudin et al., 1992). CRF is Condition Random Field, and is essentially measuring how similar the semantics of two nearby RGB pixel patches are based on how similar the RGB values are. TV and CRF losses were not included in the gradient during training.
@@
we select $\gamma=\text{med}(\left\lVert x_{i}-x_{j}\right\rVert^{2})\quad i\neq j$. We then collect results for Fidelity, TV Loss, CRF Loss, and MMD, for 4×4$\times$4 × upsampling, and display the results in table 8. We collect these results for SigLIP, DFN CLIP, and RADIO. It is clear that FeatSharp achieves the highest upsampling fidelities across the board. FeatUp produces the lowest TV and CRF losses. It achieving the lowest TV loss is intuitive given how smooth it tends to make object interiors, seen in the pca visualizations. We can see that the lower TV and CRF losses extends to FeatSharp when we apply JBU upsampling, as it achieves lower values than using bilinear upsampling for the residual pathway. The “JBU + Tiles” FeatSharp variant also does better on MDD versus “Bilinear + Tiles” across the board. It’s curious that JBU alone has the worst MMD (probably due to over-smoothing), but the best when incorporated into FeatSharp (probably owing to smoothing out the noise). We can also see that generally either “X + Tiles” FeatSharp method produces similar fidelities, aside from RADIO, where bilinear actually does do a bit better. Most likely, this is because RADIO features are themselves already fairly clean, and at some point, the structural priors of JBU actually hurt, because they’re eliminating some of the raw signal that bilinear upsampling preserves. In this case, the model always has access to the raw low-res signal with bilinear upsampling because we use an integer multiple upsampling factor, and our local attention window size is larger than this multiple. Given the totality of evidence, we choose to select “JBU + Tiles” as the default upsampling mechanism, as it’s either the best, or nearly so, across the board, and particularly, it does better with the vision models that are not able to natively change their resolution very well.
@@
Bilinear + Tiles 1.522 0.105 0.093 0.020 1.508 0.167 0.062 0.014 5.481 0.112 0.073 0.001
JBU + Tiles 1.540 0.103 0.077 0.017 1.514 0.157 0.046 0.013 5.361 0.095 0.065 0.001
@@
As can be seen in figure 14, the choice of window size has a very small impact on the resulting fidelity. However, we can also see that FeatSharp is achieving much higher fidelity scores than Bilinear and FeatUp. So, we also study what effect the attention block, and the MLP, are having on the resulting quality. We use the “Bilinear + Tile” input configuration from section B.2, and when applicable, use a window size of 5 from section B.2. We show these results in table 9. We can see that in all cases, “Attention + MLP” (e.g. the setting proposed in this work) produces the highest fidelity for both SigLIP and RADIO. However, when running the normal setup (“Short” in the table), we notice that “Attention + MLP” is usually only barely better than “MLP”, meaning that we weren’t getting much improvement from the local attention window, and that goes a long way toward explaining the relative insensitivity to the window size in figure 14. Inspired by figure 13, we notice that the longer training recipe results in much sharper images. In table 9 we study the effect of running just the MLP for the “Long” recipe. We can see that while the fidelity continues to improve, it doesn’t keep up with the “Attention + MLP” setting, demonstrating that the attention module is indeed helpful.
@@
Linear 1.444 - 1.434 - 4.134 - 4.601 -
Attention 1.437 - 1.426 - 4.017 - 4.527 -
MLP 1.444 1.514 1.433 1.514 4.151 4.668 4.599 5.128
Attention + MLP 1.452 1.544 1.442 1.525 4.216 4.957 4.656 5.361
@@ Appendix C Implementation Details
We follow the staged setup in (Heinrich et al., 2024) section 4.2, with stages 1 and 2 being exactly identical. For stage 3, in the hi-res student branch, instead of bilerp downsampling the student features to match DFN CLIP and SigLIP (RADIO-AMP baseline), we use our various upsampling methods to create hi-res feature maps which the student matches. We use our trained 3×3$\times$3 × upsamplers for the task. For FeatSharp, because we have the learned de-bias buffer which operates on the original model resolution, we also choose to apply this to the teachers in the low-res partition, as it represents the fixed bias of the teacher model, and is thus not particularly useful information.

### v2

@@ 4.3 Semantic Segmentation
Semantic segmentation has the potential to benefit from increased resolution, as it allows for label contours to be more precise, and potentially for regions to be recovered that are otherwise too small. The first setting we evaluate on is we train both FeatUp and FeatSharp at 2×2$\times$2 × and 4×4$\times$4 × upsampling, both using PHI-S. We resize the input size to be the featurizer’s native input resolution, which we call “1×1$\times$1 × Input Size”, and we also consider “2×2$\times$2 × Input Size”, where we double the input size, and feed directly to the featurizer in the case of “Baseline”, or we allow the upsampler to have higher resolution guidance while keeping the featurizer input fixed at 1×1$\times$1 × resolution. We show these results in figure 7. In most cases, both upsampling algorithms produce higher quality segmentations than the baseline, however, FeatUp is worse than the “Baseline 2×2$\times$2 ×” method for RADIOv2.5-L and ViT. In all cases, FeatSharp is superior to both FeatUp and also the baselines by significant margins. We even improve upon SOTA RADIO’s published result of 51.47 with a 2×2$\times$2 × upsampling combined with 2×2$\times$2 × input size, producing a model that attains 53.13 mIoU, a $+1.66$ mIoU improvement. RADIO itself improves with the 2×2$\times$2 × input size, but not to the same degree as with FeatSharp, with FeatSharp being $57\%$ faster. We notice that 3×3$\times$3 × upsampling is generally slightly worse than 2×2$\times$2 × or 4×4$\times$4 × for both upsamplers, but leave an investigation into why as future work. This figure also provides insight into the general ability of these foundation models to operate at resolutions that deviate from their native resolution. The CLIP family models (DFN CLIP, SigLIP, PaliGemma) are unable to benefit from this increased resolution at all, or in the case of PaliGemma, degrade with it, while the first-class-dense models like DINOv2 and RADIO natively benefit from increased resolution. Surprisingly, even though ViT is solely trained as a classification model, it also benefits from native resolution increases.
@@
4.4 Object Detection
We integrate our method, FeatUp-JBU, the baselines, as well as SAPA (Lu et al., 2022) and the preprint ReSFU (Zhou et al., 2025) into Detectron2 using the Edge [^2] https://dgcnz.github.io/edge/part2/adapting.html codebase, and probe on COCO 2017 (Lin et al., 2014). We use a [(frozen) featurizer] + [(frozen) upsampler] + ViTDet (Li et al., 2022) + DINO (Zhang et al., 2023) [^3] Not to be confused with the DINO/DINOv2 foundation models. (DETR with Improved DeNoising Anchor Boxes for End-to-End Object Detection) pipeline. We evaluate these methods on both RADIOv2.5-L (Heinrich et al., 2024) and the recently proposed SigLIP2-SO400M-512 (Tschannen et al., 2025) models. We show the results in table 1, where FeatSharp is clearly best able to improve object detection results over baseline and comparison methods, particularly for small objects, benefitting from the additional tile guidance. We also note that we were unable to use SAPA with SigLIP2 due to a CUDA configuration error in their backprop kernel.
RADIOv2.5-L
Upsampler Upsample AP
Factor * Sm Md Lg
Baseline 1 51.38 28.73 56.56 73.72
Bilinear 2 51.61 28.43 56.98 74.14
SAPA 2 41.44 15.92 45.08 69.77
ReSFU 2 49.81 26.22 55.37 73.55
FeatUp 2 46.71 21.77 52.01 72.25
FeatSharp 2 54.83 34.72 59.40 74.40
SigLIP2-SO400M-512
Baseline 1 52.66 30.31 57.94 74.31
Bilinear 2 52.69 30.19 57.84 74.16
SAPA 2 - - - -
ReSFU 2 50.84 28.45 56.18 73.69
FeatUp 2 47.42 22.87 53.17 72.80
FeatSharp 2 55.93 36.85 61.00 74.62
Table 1: COCO 2017 object detection results using Detectron2 and various upsampling methods for both RADIOv2.5-L and SigLIP2-SO400M. SAPA was unable to process this model’s input size/dimension, producing a CUDA configuration error.
@@ 4.5 Agglomerative Models
In figure 8 we qualitatively visualize the DFN CLIP adaptor features learned by the radio model. We can see that each upsampling method has a substantial impact on the resulting feature maps. The baseline method exhibits strong high-frequency artifacting starting at 768px. This is likely when RADIO “mode switches” to high-resolution, which is something that (Heinrich et al., 2024) addressed for the backbone features, but apparently still exhibit for the adaptor features. We observe that Tiling and S2 exhibit not only high-frequency noise patterns like the baseline, but also obvious grid patterns, arising from the use of tiles. More troublesome, we can see how the student learned to mimic representation switches within tiles for both Tiling and S2, where the mulch in one cell gets a different feature representation (thus color) than another, based on whether any of the dog is present in the tile view. FeatUp appears to mode switch starting at 768px into a smooth, but low-detail feature space. FeatSharp remains smooth and highly detailed as resolution increases, however, visually, it’s still possible that the features are mode switching. We show another comparison in appendix F with the SigLIP adaptor head.
Along with improvements in the adaptors, we also study the effects on the backbone features for the RADIO model. Following (Maninis et al., 2019; Lu et al., 2024) we report the MTL Gain ($\Delta_{m}$) across a suite of tasks. Unlike the prior works, instead of leveraging a single-task baseline, we opt to report the change relative to the baseline training run.
Let
@@
We show the MTL Gain results in table 2. Given that the results are relative to our baseline run, S2 and FeatSharp are the only two methods to improve, however, only FeatSharp was categorically better, leading to a +0.39% improvement across all benchmarks on average. These two methods are the only two that incorporate both low-res and hi-res features, with S2 perhaps being considered a baseline to FeatSharp, so their improvements suggest that this extra detail is indeed useful for RADIO training. We also see that our version of RADIO with FeatSharp teachers generally does better than RADIOv2.5-L (Heinrich et al., 2024), which is the current state of the art, where we improve over it on everything except for the VILA task. We report all of the raw per-task benchmark scores in tables 3, 4, 5 and 6 in the appendix.
@@ 5 Conclusion
We have presented a novel feature upsampling technique named FeatSharp that achieves higher multi-view fidelity than the current best method, FeatUp. We achieve this by joining FeatUp’s JBU upsampler with a mosaic of tiles, and then process with a single local attention block. We demonstrate its effectiveness on ADE20K semantic segmentation linear probing, where the use of FeatSharp improves over both baseline and FeatUp, even with the strongest segmenter, RADIO, which itself can handle hi-res inputs robustly. We also demonstrate our effectiveness in object detection with frozen backbone and upsampler, and see AP benefits in particular for small objects, but also medium and large. We then demonstrate the effectiveness of FeatSharp by employing it directly within RADIO training, enabling hi-res distillation targets for low-res-only teacher models. In doing so, our FeatSharp-RADIO largely improves on dense vision task benchmarks, and yields an overall improvement over our reproduction baseline, which itself improves over RADIOv2.5-L, the current state of the art. We believe this work can be useful both as a drop-in extension of existing vision systems which rely on pretrained vision encoders, as well as the newly trained FeatSharp-RADIO model with hi-res teachers, which can emulate these same models. Owing to FeatSharp-RADIO’s emulation abilities, it allows us to estimate these teacher models at arbitrary resolutions, not just integer upsampling factors as restricted in FeatSharp/FeatUp’s core training algorithm. Further, combining RADIO’s “ViTDet” (Li et al., 2022) mode with these hi-res teacher emulations allows us to achieve hi-res feature maps without fully paying the quadratic penalty in number of tokens as required by standard ViTs.
Impact Statement
This paper presents work whose goal is to advance the field of computer vision. By virtue of being a lightweight addition to existing vision models, the work aims to open up doors for higher-resolution perception tasks (e.g. segmentation, depth perception, distillation, etc.) while retaining the original model representations. As such, the ethical impacts are constrained to those of the model being upsampled. The FeatSharp training code will be released to the community.
@@ A.1 Benchmark Results
In this section, we provide detailed benchmark results used to compute the MTL aggregate metrics in table 2. We show these results in tables 3, 4, 5, and 6.
@@
Table 4: Dense and Probe3D (El Banani et al., 2024) metrics. *We report numbers for evaluation at 512px, which are found in Table A5 in RADIOv2.5 (Heinrich et al., 2024).
@@
Table 5: Pascal Context and NYUDv2 multitask learning metrics. Following the setup of MLoRE (Yang et al., 2024b) and RADIOv2.5 (Heinrich et al., 2024) with a convolutional probe. NOTE: We’re only using their harness with a conv probe, and not using their architecture.
@@ A.2 Additional Qualitative Visualizations
In figure 9, we show more PCA feature visualizations coming from our trained RADIO models. We can see that RADIO learned to mimic how tiling lacks global context, as the background-only tiles use a different feature space than those with background+content.
@@ B.2 FeatSharp Architecture
Based on figures 2 and 3, there are important degrees of freedom in the design of the system. We demonstrate in table 7 that including the de-bias module always improves the distribution matching fidelity. Here, we look at some of the other design choices:
@@
We show the results of this ablation in table 8 for both our noisiest featurizer (SigLIP), and our cleanest (RADIO) as a sanity check that we aren’t overfitting to a particular featurizer. We also note that we’re relatively agnostic to the specific base feature upsampler, allowing us to use other methods, such as ReSFU (Zhou et al., 2025), as future work.
@@
Bilinear 1.466 0.135 0.137 4.599 0.061 0.071
FeatUp 1.440 0.020 0.051 3.870 0.025 0.047
Tiles 1.261 0.897 0.237 2.713 0.357 0.094
@@
Bilinear + Tiles 1.470 0.136 0.135 4.694 0.073 0.074
FeatUp + Tiles 1.460 0.072 0.062 4.173 0.076 0.057
Tiles + Bilinear 1.337 0.812 0.241 3.202 0.336 0.098
Tiles + FeatUp 1.323 0.821 0.228 3.157 0.337 0.094
@@
Bilinear First 1.469 0.138 0.137 4.682 0.072 0.073
FeatUp First 1.473 0.063 0.065 4.202 0.064 0.057
Tiles First 1.339 0.800 0.238 3.238 0.334 0.098
Table 8: Ablation over different FeatSharp-2x configurations. Single Input means that we only supply the respective buffer to the FeatSharp module. For “Two Inputs“, we compare different low-res upsamplers in conjunction with tiling, and also the residual pathway, where the first value indicates the residual path. The FeatSharp module must integrate the information from the other value into the residual. “Three Inputs” is similar to Two, except that we only care about which buffer is the residual path, owing to the fact that there’s no intrinsic order preference in the weights for the secondary buffer(s). “TV Loss” stands for Total Variation Loss (Rudin et al., 1992). CRF is Conditional Random Field, and is essentially measuring how similar the semantics of two nearby RGB pixel patches are based on how similar the RGB values are. TV and CRF losses were not included in the gradient during training.
@@
we select $\gamma=\text{med}(\left\lVert x_{i}-x_{j}\right\rVert^{2})\quad i\neq j$. We then collect results for Fidelity, TV Loss, CRF Loss, and MMD, for 4×4$\times$4 × upsampling, and display the results in table 9. We collect these results for SigLIP, DFN CLIP, and RADIO. It is clear that FeatSharp achieves the highest upsampling fidelities across the board. FeatUp produces the lowest TV and CRF losses. It achieving the lowest TV loss is intuitive given how smooth it tends to make object interiors, seen in the pca visualizations. We can see that the lower TV and CRF losses extends to FeatSharp when we apply JBU upsampling, as it achieves lower values than using bilinear upsampling for the residual pathway. The “JBU + Tiles” FeatSharp variant also does better on MMD versus “Bilinear + Tiles” across the board. It’s curious that JBU alone has the worst MMD (probably due to over-smoothing), but the best when incorporated into FeatSharp (probably owing to smoothing out the noise). We can also see that generally either “X + Tiles” FeatSharp method produces similar fidelities, aside from RADIO, where bilinear actually does do a bit better. Most likely, this is because RADIO features are themselves already fairly clean, and at some point, the structural priors of JBU actually hurt, because they’re eliminating some of the raw signal that bilinear upsampling preserves. In this case, the model always has access to the raw low-res signal with bilinear upsampling because we use an integer multiple upsampling factor, and our local attention window size is larger than this multiple. Given the totality of evidence, we choose to select “JBU + Tiles” as the default upsampling mechanism, as it’s either the best, or nearly so, across the board, and particularly, it does better with the vision models that are not able to natively change their resolution very well. We also note that newer methods, as they emerge, could serve as better core upsampler modules than bilinear/JBU, and can be trivially swapped in.
@@
Bilinear + Tiles 1.522 0.105 0.093 0.020 1.484 0.167 0.062 0.014 5.921 0.112 0.073 0.001
JBU + Tiles 1.580 0.103 0.077 0.017 1.493 0.157 0.046 0.013 5.898 0.095 0.065 0.001
@@
As can be seen in figure 13, the choice of window size has a very small impact on the resulting fidelity. However, we can also see that FeatSharp is achieving much higher fidelity scores than Bilinear and FeatUp. So, we also study what effect the attention block, and the MLP, are having on the resulting quality. We use the “Bilinear + Tile” input configuration from section B.2, and when applicable, use a window size of 5 from section B.2. We show these results in table 10. We notice that it’s not until the long recipe where at inclusion of attention is helpful, and that goes a long way toward explaining the relative insensitivity to the window size in figure 13. Inspired by figure 12, we notice that the longer training recipe results in much sharper images. In table 10 we study the effect of running just the MLP for the “Long” recipe. We can see that while the fidelity continues to improve, it doesn’t keep up with the “Attention + MLP” setting, demonstrating that the attention module is indeed helpful.
@@
Linear 1.521 1.581 1.508 1.566 4.702 5.397 4.926 5.701
Attention 1.505 1.566 1.500 1.560 4.410 5.253 4.656 5.446
MLP 1.522 1.581 1.506 1.566 4.741 5.397 4.934 5.707
Attention + MLP 1.513 1.584 1.502 1.568 4.668 5.502 4.849 5.711
@@ Appendix C Implementation Details
We follow the staged setup in (Heinrich et al., 2024) section 4.2, with stages 1 and 2 being exactly identical. For stage 3, in the hi-res student branch, instead of bilerp downsampling the student features to match DFN CLIP and SigLIP (RADIOv2.5 baseline), we use our various upsampling methods to create hi-res feature maps which the student matches. We use our trained 3×3$\times$3 × upsamplers for the task, as they’re the smallest factor that produces feature maps larger than RADIO’s hi-res partition. For FeatSharp, because we have the learned de-bias buffer which operates on the original model resolution, we also choose to apply this to the teachers in the low-res partition, as it represents the fixed bias of the teacher model, and is thus not particularly useful information.